# Normal liquid 3He studied by Path Integral Monte Carlo with a parametrized partition function

### v1

@@ IV Finite-size effects
In order to study how these limitations would affect the extrapolation to the fermion ($\xi=-1$) case, we can use the recursion formula for the canonical partition function developed in Ref. 19 for bosons and fermions, generalized to the case of arbitrary $\xi$ as
@@
where $N$ is the number of particles and $S(l)=\sum_{i,j,k}e^{l\beta\epsilon(q_{i},q_{j},q_{k})}$[20, 21] In the following, we will set $N=64$, which is the same number of particles used in the PIMC simulations of 3He.
We show in Fig. 3(a) and 3(c) the $E_{T}(\xi)$ curves for temperatures $T=6$ K and $T=8$ K, respectively. One can see that in a finite-size system the non-analytic behavior across the BEC transition is smoothed, as evidenced by the behavior of the derivative $\partial{E_{T}}/{\partial\xi}$ (short-dashed blue lines). Additionally, the energy per particle in the fermion case is also slightly increased with respect to the bulk value (long-dashed lines).
Figure 3: (a) Comparison of the energy and its derivative with respect to $\xi$ for $T=6$ K for the bulk system (black points and black short-dashed line) and the finite size system (blue points and blue dashed-line). The long-dashed lines are located at the Fermi level of the bulk (black) and finite system (blue). (b) Comparison of $\xi_{E}(T)$ for $E-E_{F}=2.512$ K for the bulk system (black circles) versus the finite system (blue triangles). This value for the energy is chosen at the $\xi_{\mathrm{c}}$ of the bulk system in (a). (c) Same as (a) but for $T=8$ K. (d) same as (b) but at the critical point of (c), corresponding to an energy of $E-E_{F}=3.477$ K.
The finite-size smoothing of the $E_{T}(\xi)$ curves is mirrored into the $\xi_{E}(T)$ curves, as can be seen in Figs. 3(b) and 3(d) (blue triangles). The inflection due to BEC is much less evident, and finite-size effects generally flatten the $\xi_{E}(T)$ curve in its proximity. Nevertheless, the energy in the fermion limit ($\xi=-1$) is changed only slightly. In this case, the critical value of $\xi$ turns out to be very close to 0, preventing a a good fit with the functional form of Eq. (14). Therefore, we found it necessary to slightly change the way of extrapolating results to the fermion case. We notice that the behavior of $\xi_{E}(T)$ close to $\xi=0$ is effectively linear, and hence we adopt a fitting function of the form:
@@
In Figures 3(b) and 3(d) the red dashed line represents the fit using Eq. (16) taking into account only $\xi_{E}\geq 0$ finite size points, while the green continuous curve is the fit taking into account the points for which $\xi_{E}\geq-1$. Notice that the linear extrapolation introduces a slight uncertainty in the order of $0.1$ K in the position of the temperature where $\xi=-1$ is crossed.
V PIMC results for 3He
Figure 4: 3He results. (a) Energy as a function of the $\xi$ parameter for temperatures $T$ in the range of ($1.294$, $3.667$) K. (b-d) $\xi_{E}(T)$ for four different energies per particle. We report the values of the fit using Eq. (16) (red lines), the linear fit passing through the points for $\xi\sim 0$ (dark-blue dashed lines) and the cubic fit using Eq. (6) (dark-green dashed lines). (e) Total energy per particle as a function of temperature. Black line corresponds to the experimental data from Ref. [22]; red points are the one extrapolated to $\xi=-1$, light-gray triangular points are the $\xi=0$ (Boltzmannons) energies while dark-gray square points are the numerical results obtained using PIMC plus fixed-node approximation in Ref. [8].
@@ V PIMC results for 3He
In the case of 3He simulations, we used the PIMC approach [23, 15] to sample the partition function and the worm algorithm to sample permutations. [24] Details of PIMC algorithm are presented elsewhere. [25, 26] The only modification to the original set of MC moves,[25] is the one including the $\xi$ factor of Eq. (2) in the weight factor of the swap move, the one responsible for the exchange permutation sampling. We considered a system of $N=64$ atoms at a density $\rho=0.016355$ atoms/Å3, using periodic boundary conditions. We employed the most recent ab-initio two-body potential to treat the interaction between 3He atoms [27] and we evaluated the propagator in the pair-product approximation [28] with fixed $\tau=\frac{\beta}{P}=\frac{1}{44}$ K-1, where $P$ is the number of beads (imaginary-time slices) of the PIMC approach. For each temperature, $P$ is thus chosen as $P=\text{int}\left[\frac{1}{\tau T}\right]$, where $\mathrm{int}(x)$ is the closest integer number to $x$. Energies are estimated using the virial estimator [15] and then are corrected perturbatively by adding the three-body potential contribution. [29] Results for 3He are shown in Fig. 4. Panel (a) reports the energies obtained as a function of the $\xi$ parameter at temperatures ranging from $1.294$ K to $3.667$ K. The points obtained from PIMC calculations are interpolated using B-splines. [30] The behaviour of the $E_{T}(\xi)$ closely resembles the finite size calculations using the helium-like dispersion of Eq. (13), with a $T$-dependent smoothed inflection point where the superfluid transition occurs. In Figs. 4(b-d) we show three $\xi_{E}(T)$ functions, estimated by cutting the curves in Fig. 4(a) at constant energies. Blue points represent numerical data, while the red lines are the fits using Eq. (16). For the sake of reproducibility, the values of the parameters are reported in those figures. The dark-green dashed lines represent fits using a simple cubic polynomial, as used in the original papers on this method; [12] Notice that in some cases (panels (b) and (c)), this fitting function does not intersect the $\xi=-1$ line, which leads to an unphysical result: it would imply that no temperature corresponds to the assumed value of the energy per particle.
@@
Finally, we show in Fig. 4(e) and Table 1 the results for the energy per particle obtained by extrapolation of the fitting function in Eq. (16). The experimental results from Ref. 22 are in remarkably close agreement with our findings. In Fig. 4(e), we also report the energy per particle without exchange effects (Boltzmannons, that is $\xi=0$) and the results obtained using PIMC in conjunction with a fixed-node constraint: [8] the former strongly underestimate the fermion energies, while the latter are closed to the experimental value, but display a more scattered behavior.
$E/N$ [K] $T$ [K] Extrapol. $\xi=-1$ $T$ [K] Boltzmann. $\xi=0$
$-2.445$ $0.149$ $1.294$
$-2.310$ $0.593$ $1.467$
$-2.224$ $0.714$ $1.571$
$-2.018$ $1.189$ $1.833$
$-1.880$ $1.470$ $2.000$
$-1.710$ $1.694$ $2.200$
$-1.480$ $2.081$ $2.444$
@@
Table 1: Summary of the numerical energies and temperatures reported in Fig. 4(e). All the energies are given in Kelvin units.
VI Conclusions
@@
In this paper, we presented simulations of normal liquid 3He where the fermionic nature of the system was taken into account using a recently developed approach based on a $\xi$-parametrized partition function that continuously interpolates between bosonic ($\xi=1$) and fermionic ($\xi=-1$) statistics. We found that the presence of the superfluid transition in the $\xi>0$ sector, where calculations were not affected by the sign problem, induced significant non-analytic behavior in the function $E(T,\xi)$ that required a tailored approach to extrapolate to the fermion case. Our simulations yielded very good agreement with experimental data for the energy of particles. The approach developed here may find useful application in other strongly-interacting fermionic systems, such as those relevant to nuclear physics.
@@ Acknowledgements
TM and GG thank Gabriele Spada and Stefano Giorgini for useful discussions and for the details of the PIMC+worm algorithm.

### v2

@@ III.1 Finite-size effects
In the case of PIMC simulations, the results of energy calculations are generally influenced by finite-size effects. In order to study how these limitations would affect the extrapolation to the fermion ($\xi=-1$) case, we can use the recursion formula for $N$ identical particles in the canonical ensemble developed in Ref. 21 for bosons and fermions, generalized to the case of arbitrary $\xi$ as
@@
where $S(l)=\sum_{i,j,k}\text{exp}\left[l\beta\epsilon(q_{i},q_{j},q_{k})\right]$[22, 23] For a system composed of particles with different spins, the discrete partition function is straightforwardly generalized as $Z(N,\xi)=\prod_{\sigma}Z(N_{\sigma,\xi})$. In the following, we will set $N\equiv N_{\uparrow}+N_{\downarrow}=64$, which is the same number of particles used in the PIMC simulations of 3He.
We show in Fig. 4(a) and 4(c) the $E_{T}(\xi)$ curves for $T=6$ K, in the fully polarized case ($P=1$) and unpolarized ($P=0$) case, respectively. One can see that in a finite-size system the non-analytic behavior across the BEC transition is smoothed, as evidenced by the behavior of the derivative $\partial{E_{T}}/{\partial\xi}$ (blue lines).
The finite-size smoothing of the $E_{T}(\xi)$ curves is mirrored into the $\xi_{E}(T)$ curves, as can be seen in Figs. 4(b) and 4(d) (blue triangles). The inflection due to BEC is much less evident, and finite-size effects generally flatten the $\xi_{E}(T)$ curve in its proximity. Nevertheless, the energy in the fermion limit ($\xi=-1$) is changed only slightly. In this case, the critical value of $\xi$ turns out to be very close to 0, preventing a a good fit with the functional form of Eq. (17). Therefore, we found it necessary to slightly change the way of extrapolating results to the fermion case. We notice that the behavior of $\xi_{E}(T)$ close to $\xi=0$ is effectively linear, and hence we adopt a fitting function of the form:
@@
In Figures 4(b) and 4(d) the red dashed line represents the fit using Eq. (19) taking into account only $\xi_{E}\geq 0$ finite size points, while the green continuous curve is the fit taking into account the points for which $\xi_{E}\geq-1$. Note that the linear extrapolation introduces a slight uncertainty, $\delta T_{\mathrm{fs}}\eqsim 0.15$ K in the position of the temperature at which $\xi=-1$ is crossed.
Figure 5: PIMC results for 3He. Energy as a function of the $\xi$ parameter for temperatures $T$ in the range of ($1.294$, $3.667$) K in the fully polarized (panel (a)) and unpolarized (panel (b)) cases.
Figure 6: Plot of $\xi_{E}(T)$ in the 3He case from PIMC simulations for two different energies per particle $E/N=-2.445$ K (panels (a) and (b)) and $E/N=-2.018$ K (panels (c) and (d)). Panels (a) and (c) are obtained in the fully polarized case while (b) and (d) in the unpolarized one. We report the values of the fit using Eq. (19) (red lines), the linear fit passing through the points for $\xi\sim 0$ (orange long-dashed lines) and the cubic fit using Eq. (7) (dark-green dashed lines). The value of $T_{\mathrm{c}}$ has been estimated using the procedure reported in the main text.
IV PIMC results for 3He
@@
In the case of 3He simulations, we used the PIMC approach [24, 16] to sample the partition function and the worm algorithm to sample permutations. [25] Details of PIMC algorithm are presented elsewhere. [26, 27] The only modification to the original set of MC moves,[26] is the one including the $\xi$ factor of Eq. (2) in the weight factor of the swap move, the one responsible for the exchange permutation sampling. We considered a system of $N=64$ atoms at a density $\rho=0.016355$ atoms/Å3, using periodic boundary conditions. We employed the most recent ab-initio two-body potential to treat the interaction between 3He atoms [28] and we evaluated the propagator in the pair-product approximation [29] with fixed $\tau=\frac{\beta}{P}=\frac{1}{44}$ K-1, where $P$ is the number of beads (imaginary-time slices) of the PIMC approach. For each temperature, $P$ is thus chosen as $P=\text{int}\left[\frac{1}{\tau T}\right]$, where $\mathrm{int}(x)$ is the closest integer number to $x$. Energies are estimated using the virial estimator [16] and then are corrected perturbatively by adding the three-body potential contribution. [30] Results of the $E_{T}(\xi)$ for 3He are shown in Fig. 5. Panel (a) reports the energies obtained as a function of the $\xi$ parameter at temperatures ranging from $1.294$ K to $3.667$ K for $P=1$ while panel (b) the same $E_{T}(\xi)$ but for $P=0$. The points obtained from PIMC calculations are interpolated using B-splines. [31] The behaviour of the $E_{T}(\xi)$ closely resembles the finite-size calculations using the helium-like dispersion of Eq. (16), with a $T$-dependent smoothed inflection point where the superfluid transition occurs.
Similar to the independent-particle system discussed above, both the unpolarized and polarized systems converge to the same value of the energy for $\xi\to 0$ and the inflection points of $E_{T}(\xi)$ are shifted towards larger $\xi$-values in the unpolarized case. This is a consequence of the fact that the unpolarized system is ”less bosonic” as only half of the particles participate in permutational exchanges. Consequently, the energy remains closer to $E_{T}(\xi=0)$ for larger values of $\xi$.
Figure 7: Total energy per particle as a function of temperature from PIMC simulations. Black line corresponds to the experimental data from Ref. [32]; red diamonds are the results extrapolated to $\xi=-1$ in the fully polarized case ($P=1$), blue circles are the one extrapolated to $\xi=-1$ in the unpolarized case ($P=0$), light-gray triangular points are the $\xi=0$ (Boltzmannons) energies while dark-gray square points are the numerical results obtained using PIMC plus fixed-node approximation in Ref. [8]. Details of error bar computation are provided in the text.
We show in Fig. 6 some $\xi_{E}(T)$ functions, estimated by cutting the curves in Fig. 5(a) (see Figs. 6(a) and (c)) and in Fig. 5(b) (see Figs. 6(b) and (d)) at constant energies. Blue points represent numerical data, while the red lines are the fits using Eq. (19). For the sake of reproducibility, the values of the parameters are reported in those figures. The dark-green dashed lines represent fits using a simple cubic polynomial, as used in the original papers on this method. [12] Notice that in all the reported cases this fitting function does not intersect the $\xi=-1$ line, which leads to an unphysical result: it would imply that no temperature corresponds to the assumed value of the energy per particle. Fitting Eq. (19) to the numerical data for a given energy $E$, turns our to be more difficult for the unpolarized case (panels (b) and (d)) than for the polarized case since the $T_{\mathrm{c}}$ is more difficult to estimate from the available data. We estimated it using the following procedure: we fitted the data with a function which is linear below a certain $T_{0}$ and quadratic above. For each value of $T_{0}$ we computed the $\chi$-squared of the optimal fit, $\chi^{2}(T_{0})$, and we chose $T_{c}$ as the position of the minimum of $\chi^{2}(T_{0})$. We emphasize that this procedure would be much smoother for larger number of particles $N$.
Finally, we show in Fig. 7 and Table 1 the results for the energy per particle obtained by extrapolation of the fitting function in Eq. (19). We also report an estimate of the uncertainty of our results, coming from the uncertainty in the location of the solution of $\xi_{E}(T)=-1$. This contribution to the uncertainty is larger in the fully polarized case than in the unpolarized one, as in the independent-particle model (see Fig. 3 above) and has been estimated by varying the $T_{\mathrm{c}}$ in Eq. (19) by $0.2$ K and re-extrapolating. Specifically, this variation, denoted as $\delta T_{\mathrm{c}}$, represents the main source of error in determining $E(T)/N$, being larger than the uncertainties arising from the PIMC simulations and the interpolation of $E_{T}(\xi)$ data points at $\xi\geq 0$. Considering also the systematic uncertainty estimated from the independent-particle model in the finite size case, that is the contribution $\delta T_{\mathrm{fs}}$ discussed in sec. III.1, the error bar $\delta T$ shown in Fig. 7 was computed as $\delta T=\sqrt{\delta T_{\mathrm{c}}^{2}+\delta T_{\mathrm{fs}}^{2}}$.
@@
As expected from the independent particle model and from previous results, [33] we found that the energy of the spin-unpolarized system is lower than that of the fully polarized one by $\approx 0.2$ K. The same figure also shows the experimental results from Ref. 32, which have been obtained by integrating the specific heat curve and then rigidly shifted by $-2.4788$ K [33] to match the experimental value at $T=0$. Our results show a fairly good agreement with experimental data considering the uncertainties of our calculation, although we notice a systematic underestimation of the energy per particle by $\approx 0.2$ K. We also report in Fig. 7 the energy per particle without exchange effects (Boltzmannons, that is $\xi=0$, light-grey trinagles) and the results obtained using PIMC in conjunction with a fixed-node constraint (dark-grey squares): [8] the former strongly underestimate the fermion energies, while the latter are closed to the experimental value, but display a more scattered behavior.
$E/N$ [K] $T$ [K], $\xi=-1$ $T$ [K], $\xi=-1$ $T$ [K], $\xi=0$
$P=1$ $P=0$
$-2.445$ $0.19\pm 0.30$ $0.57\pm 0.22$ $1.29$
$-2.400$ $0.40\pm 0.38$ $0.66\pm 0.21$ $1.33$
$-2.310$ $0.60\pm 0.29$ $0.94\pm 0.21$ $1.45$
$-2.224$ $0.71\pm 0.21$ $1.19\pm 0.23$ $1.56$
$-2.018$ $1.19\pm 0.20$ $1.46\pm 0.19$ $1.82$
$-1.880$ $1.47\pm 0.21$ $1.64\pm 0.20$ $1.99$
$-1.710$ $1.69\pm 0.18$ $1.92\pm 0.18$ $2.18$
$-1.480$ $2.08\pm 0.24$ $2.20\pm 0.17$ $2.44$
@@
Table 1: Summary of the numerical energies and temperatures reported in Fig. 7. All the energies are given in Kelvin units.
V Conclusions and perspectives
@@
We presented simulations of normal liquid 3He where the fermionic nature of the system was taken into account using a recently developed approach based on a $\xi$-parametrized partition function that continuously interpolates between bosonic ($\xi=1$) and fermionic ($\xi=-1$) statistics. By studying an independent-particle model where results are numerically exact, we found that the presence of the superfluid transition in the $\xi>0$ sector, where calculations were not affected by the sign problem, induced significant non-analytic behavior in the function $E(T,\xi)$ that required a tailored approach to extrapolate to the fermion case.
Our simulations, performed for polarized and unpolarized liquid 3He, yielded good agreement with experimental data for the energy of particles, although the $P=0$ results underestimate systematically the experimental curve by $\sim 0.2$ K. While the aim of the work was the extension of the formalism previously developed in Ref. 12, the approach developed here may find useful application in other strongly-interacting fermionic systems, such as those relevant to nuclear physics.
@@ Acknowledgements
TM and GG thank Gabriele Spada and Stefano Giorgini for useful discussions and for the details of the PIMC+worm algorithm. We also extend our gratitude to an anonymous reviewer for their valuable feedback and constructive comments.
Appendix A Independent-particle model: Polarized vs unpolarized energies
While the energy of the independent-particle model has been reported in Figs. 2(a) and  2(c) for the polarized and unpolarized cases respectively, here we directly compare those energies in figure 8.
Figure 8: Independent-particle model. Energies as a function of the $\xi$ parameter for the dispersion in Eq. (16) of the main text and for $P=1$ (continuous lines) and $P=0$ (circles) from Eqs. (10) and (11) of the manuscript for $T=6$ (red color), $T=8$ (blue color) and $T=10$ K (black color).
The direct comparison is useful to highlight the different behaviour of the $P=0$ (points) and $P=1$ (continuous lines) energies in the range of $\xi$ including all the physical values, i.e. $\xi=-1,0,1$. From Fig. 8 one can observe the flattening of the energy dispersion around the Boltzmannon energy ($\xi=0$) induced by the condition $P=0$. In the latter case, the system consists of two distinct types of particles with different spins. As a result, for a fixed total density, the system is less fermionic and less bosonic since only half of the particles are involved in permutational exchanges. Therefore, for a spin-independent Hamiltonian, the fermionic system always favours the unpolarized state having lower energy.
Appendix B Independent-particle model: Free dispersion
By using the same dispersion of Eq. (16) with $\alpha_{1}=0$, we solved numerically Eqs. (10) and (11) fixing again the density at $mN/\Omega=0.081$ g/cm3. Results are shown in Fig. 9 for the fully polarized system. In panel (a) we report the $E_{T}(\xi)$ dispersions for temperatures ranging from $1$ to $10$ K. Red stars indicate the critical $\xi$ points at which the system undergoes the BEC transition.
Figure 9: Independent–particle model. (a) Energies as a function of the $\xi$ parameter for the free dispersion (obtained by setting $\alpha_{1}=0$ in Eq. (16) of the manuscript) and for $P=1$ (fully polarized case) from Eqs. (10) and (11) for temperatures between $1$ K and $10$ K. Red stars indicate the energies obtained at $\xi_{\mathrm{c}}$ in Eq. (12). (b) Behaviour of $\xi_{E}(T)$ for a few selected energies for $P=1$. The latter are chosen at the critical points for $6,8$, and $10$ K. Notice the absence of slope change of $\xi_{E}(T)$ at these temperatures, at variance with Fig. 2(b).
We notice that at variance with the behaviour of $E_{T}(\xi)$ in Fig. 2(a), here the inflection point is much smoother due to the different single-particle dispersion.
This is reflected in panel (b) in the shape of the $\xi_{E}(T)$, which follows a parabolic behaviour as predicted in Ref. [12]. In this case, Equation 17 clearly represents an overfit to $\xi_{E}(T)$ as the critical temperature is not evident. Nevertheless, Eq. (17) automatically embodies the possible inflection point that can be driven by superfluid effects in strongly correlated systems.